OpenStudy (anonymous):

PLEASE help assignment due in half an hour!!!!! A baseball team plays in he stadium that holds 68000 spectators. With the ticket price at 9 the average attendence has been 29000. When the price dropped to 8, the averege attendence rose to 34000. Find the demand function p(x), where x is the number of the spectators. (assume p(x) is linear)

OpenStudy (anonymous):

Only final answer is needed. No explanation.

OpenStudy (anonymous):

56

OpenStudy (anonymous):

I need a demand function...

OpenStudy (lilg132):

(x1, y1) = (9,29k) (x2, y2) = (8,34k) y - 34k = [(34 - 29)/(8-9)](x-9) y - 34 = -5(x-9) y - 34 = -5x + 45 y = -5x + 79 please dont forget to give a medal :)